### v2

# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

Abstract
We elucidate the generic bifurcation behavior of local and global order in the short-range interacting nonreciprocal Ising model evolving under Glauber dynamics. We show that a critical magnitude of nearest-neighbor correlations within the respective lattices controls the emergence of coherent oscillations of global order as a result of frustration. Local order is maintained during these oscillations, implying non-trivial spatio-temporal correlations. Long-lived states emerge in the strong-interaction regime. The residence time in either of these states eventually diverges, giving rise to ordered non-equilibrium steady states via a saddle-node-infinite-period bifurcation. Our work provides a comprehensive microscopic understanding of collective phenomena in nonreciprocal matter without conservation laws.
The last decade saw a surge of interest in multicomponent systems with nonreciprocal interactions. Nonreciprocity is commonly found in active and non-equilibrium systems [1, 2, 3, 4, 5], such as predator-prey [6, 7] and many-body lattice [8, 9, 10, 11, 12, 13] systems, spin glasses [14, 15, 16], active solids with odd elasticity [17], cell-cell interactions [18], optically trapped nanoparticles [19, 20], complex plasma dynamics [21, 22, 23], mixtures of passive and active particles [24] as well as chemically active interacting particles [25, 26, 27, 28, 29], primitive metabolic cycles [30], and non-Hermitian quantum systems [31, 32, 33]. The importance of nonreciprocal interactions is also widely recognized in the social sciences [34, 35], neuroscience [36, 37], and robotics [38]. However, the underlying mechanisms and implications of nonreciprocity can vary substantially among these disciplines.
At the microscopic level, nonreciprocal interactions are closely linked to the violation of Newton’s third law and broken detailed-balance [21, 2]. On the collective level, nonreciprocally interacting systems can resist coarsening and instead self-organize into dynamic states with unique temporal and spatial patterns, including traveling and oscillatory states [39, 40, 41, 42, 43, 44, 45, 46]. Phenomenogically, such systems are typically described by a nonvariational coupling of Cahn-Hilliard models or Allen-Cahn models (i.e., the respective models B and A in [47]).
An important aspect of nonreciprocal systems is the presence of conservation laws. In the case of two nonvariationally coupled fields one may have no [48, 8, 16, 49], one [50, 51, 52], or two [2, 24, 40, 41, 46, 43, 53] conservation laws. Nonreciprocal systems without conservation laws are particularly relevant in the realm of chemical reaction networks, where it was shown that kinetic asymmetries between enzymes may lead to nonreciprocal interactions, resulting in spatio-temporal oscillations [28].
There is growing interest in a bottom-up understanding of non-conserved nonreciprocal systems, so far limited to phenomenological and mean-field theory [48, 8, 16]. These works revealed intriguing phenomena, such as large-scale Hopf instabilities [48, 8, 16], saddle-node bifurcations [49, 48], and hidden collective oscillations [16].
Here, we go beyond mean-field reasoning and develop a thermodynamically consistent description of nonreciprocal binary systems without conservation laws, starting from a pair of coupled Ising models, which serves as a minimal model of a many-body system with short-range interactions. We incorporate nearest-neighbor correlations and consider both, global and local order. We explain why a critical magnitude of nearest-neighbor correlations controls the symmetry-breaking transition in the global order, in turn bounding the onset of coherent oscillations. We elucidate how the bifurcation behavior depends on inter-particle interactions, and highlight stark differences in the spatio-temporal dynamics of all-to-all (mean-field) versus short-range-interacting systems; the square and Bethe lattices display equivalent behavior that is strikingly different from the all-to-all lattice.
Fundamentals.—Consider a pair of lattices denoted by $\mu=a,b$, as shown in Fig. 1a, each having a coordination number $z$ and periodic boundary conditions. On each lattice there are $N$ spins that can assume two states $\sigma^{\mu}_{i}=\pm 1$, with $i\in\{1,...,N\}$ enumerating the spin’s location. Each spin interacts with its $z$ nearest-neighbors on the same lattice and the spin at the equivalent position on the opposing lattice. The local interaction  [^1] The local interaction is not equal to the total energy of the system, since the interaction energy for the two lattices $a$ and $b$ is different. energy of spin $i$ on lattice $\mu$ can be written as (see also [48])
$E^{\mu}_{i}=-J_{\mu}\sigma^{\mu}_{i}\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}-K_{\mu}\sigma^{a}_{i}\sigma^{b}_{i},\$ (1)
where $\langle i|j\rangle$ denotes a sum over nearest-neighbors $j$ of spin $i$. Throughout, we will express energies in units of $k_{\rm B}T$, where $T$ is the temperature of the heat bath. The parameters $J_{\mu}$ and $K_{\mu}$ denote the coupling within lattice $\mu$ and of the spins in $\mu$ with those on the opposing lattice, respectively. When $K_{a}\neq K_{b}$, equivalent spins on the opposing lattices interact nonreciprocally.
Figure 1: (a) Schematic of two Ising lattices $a$ and $b$ with coordination number $z{=}4$ and cross-coupling $(K_{a},K_{b})$. For $K_{a}\neq K_{b}$ the cross-coupling is nonreciprocal. (b) Eigenvalues $\lambda_{\pm}$ of the linear stability matrix (see Eq. (8)) as a function of the steady state local order $q_{s}$ for $K_{a}=-K_{b}=0.1$. When $q_{s}\geq q_{\rm crit}\equiv 1/(z-1)$ (black vertical line) the real parts of $\lambda_{\pm}$ are non-negative, resulting in coherent oscillations. (c-f) Temporal evolution of the magnetization (c; see Eq. (2)), (d) local order (see Eqs. (3)-(4)), and (e,f) local correlations $\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t)$ for $K_{a}=-K_{b}=K=0.1$ and $J_{a}=J_{b}=0.4$. Black dashed lines in (c,d) are obtained with Monte-Carlo simulations on the Bethe lattice (see [55] for details). In all panels we consider $z=4$.
We consider single spin-flip Glauber dynamics [56]. Let $P(\boldsymbol{\sigma};t)$ be the probability at time $t$ to find the system in state $\boldsymbol{\sigma}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ which is governed by the master equation $dP(\boldsymbol{\sigma};t)/dt=\sum_{\mu}\sum_{i}[w^{\mu}_{i}(-\sigma^{\mu}_{i})P(\boldsymbol{\sigma}^{\prime}_{\mu,i};t){-}w^{\mu}_{i}(\sigma^{\mu}_{i})P(\boldsymbol{\sigma};t)]$, where $\boldsymbol{\sigma}^{\prime}_{\mu,i}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,-\sigma^{\mu}_{i},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ is a state which differs from state $\boldsymbol{\sigma}$ by one spin flip. The transition rates $w^{\mu}_{i}(\sigma^{\mu}_{i})$ to flip a spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$ are uniquely specified by limiting the interactions to nearest neighbors, imposing isotropy in position space, and requiring that for $K_{a}=K_{b}$ the transition rates obey detailed-balance. These physical restrictions then lead to the general result $w^{\mu}_{i}(\sigma^{\mu}_{i}){=}\left[1{-}\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\right]/2\tau$ [8, 48, 16], where $\Delta E^{\mu}_{i}=-2E^{\mu}_{i}$ is the change in energy on the $\mu=a\lor b$ lattice after spin conversion $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau$ is an intrinsic timescale to attempt a single spin-flip.
We are interested in the temporal dynamics of global and local order parameters averaged over all spins. The magnetization or global order [57] is given by
$m^{\mu}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{\mu}_{i}(t)\rangle,$ (2)
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Note that $m^{\mu}(t)\in[-1,1]$. The three local order parameters [57] are
$q^{\mu\mu}(t)$ $\equiv\frac{2}{zN}\sum_{i}\sum_{\langle i|j\rangle}\langle\sigma_{i}^{\mu}(t)\sigma_{j}^{\mu}(t)\rangle,$ (3)
$q^{ab}(t)$ $\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma_{i}^{a}(t)\sigma_{i}^{b}(t)\rangle,$ (4)
with correlations $\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t)$. We distinguish between the local order within and between the lattices. The alignment of spin pairs within lattice $\mu$ is quantified by $q^{\mu\mu}(t)$, and $q^{ab}(t)$ measures the alignment of equivalent spins on opposing lattices. The normalization in Eq. (3) arises because $zN/2$ is the number of nearest neighbor pairs in a periodic lattice with coordination number $z$, which renders $q^{\mu\mu}(t),q^{ab}(t)\in[-1,1]$.
Evolution equations.—Our first main result is an exact set of coupled differential equations for the order parameters in the thermodynamic limit $N\rightarrow\infty$, which reads (see [55] for a detailed derivation)
$\tau\frac{dm^{\mu}(t)}{dt}{+}m^{\mu}(t)$ ${=}\sum\nolimits_{l,n}\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$
$\tau\frac{dq^{\mu\mu}(t)}{dt}{+}2q^{\mu\mu}(t)$ ${=}\frac{2}{z}\!\sum\nolimits_{l,n}\!(2l{-}z)\mathcal{P}^{\mu}_{l,n}(t)\!\tanh(U^{\mu}_{l,n}),$ (5)
$\tau\frac{dq^{ab}(t)}{dt}{+}2q^{ab}(t)$ ${=}\!\sum\nolimits_{\mu}\sum\nolimits_{l,n}(2n{-}1)\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$
where $\sum_{l,n}\equiv\sum_{l=0}^{z}\sum_{n=0}^{1}$ is a sum over all possible nearest neighbor spin values on the same ($l=0,\ldots,z$) and opposing ($n=0,1$) lattice, $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$, and $\mathcal{P}^{\mu}_{l,n}(t)\in[0,1]$ is the time-dependent probability to select an up or down spin which has $l$ neighboring up spins on the same lattice, and $n$ neighboring up spins on the opposing lattice. The probability is normalized as $\sum_{l,n}\mathcal{P}^{\mu}_{l,n}(t){=}1$. Equations (5) are not yet closed; evaluating $\mathcal{P}^{\mu}_{l,n}(t)$ for an arbitrary lattice is a daunting combinatorial task, as it depends on microscopic details and therefore on an infinite hierarchy of order parameters.
An accurate closed-form expression for $\mathcal{P}^{\mu}_{l,n}(t)$ can be obtained with the Bethe-Guggenheim (BG) approximation, where we assume perfect mixing of nearest neighbor spin pairs. Such an approximation is exact on loop-less lattices like the Bethe lattice [58], or large random graphs with fixed coordination number [59, 60]. As we show below, the spatio-temporal dynamics on these lattices captures the behavior on the square lattice. Note that a theory of the nonreciprocal Ising model on the fully-connected mean-field lattice has been constructed in [48], which we also report in Appendix A.
We split $\mathcal{P}^{\mu}_{l,n}(t)$ in “up” and “down” spin contributions $\mathcal{P}^{\mu}_{l,n}(t)=\mathcal{P}^{\mu+}_{l,n}(t)+\mathcal{P}^{\mu-}_{l,n}(t)$, where, e.g., $\mathcal{P}^{\mu+}_{l,n}(t)\in[0,1]$ is the probability of flipping an up spin with $l$ up neighbors on the same and $n$ up neighbors on the opposing lattice, respectively. These probabilities read  [^2] See Supplementary Material at […] (omitting the explicit $t$-dependence on the r.h.s.)
$\mathcal{P}^{a\pm}_{l,n}(t){=}\frac{C^{z}_{l}(1{\pm}2m^{a}{+}q^{aa})^{\delta^{\pm}_{l}}(1{\pm}m^{a}{-}m^{b}{\mp}q^{ab})^{1-n}}{(1{\pm}m^{a})^{z}(1{-}q^{aa})^{-\delta^{\mp}_{l}}(1{\pm}m^{a}{+}m^{b}{\pm}q^{ab})^{-n}},$ (6)
where $C^{z}_{l}{\equiv}\binom{z}{l}/2^{z+2}$, $\delta^{+}_{l}{=}l$, and $\delta^{-}_{l}{=}z-l$. The expression for $\mathcal{P}^{b\pm}_{l,n}(t)$ follows from Eq. (6) by replacing $m^{a}{\rightarrow}m^{b}$ and $q^{aa}{\rightarrow}q^{bb}$. Inserting Eq. (6) into Eqs. (5) yields a closed system of five coupled nonlinear differential equations.
Linear stability analysis.—We focus on the symmetric nonreciprocal setting $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, also known as the perfectly nonreciprocal [15] setting. We start with the linear stability analysis of Eqs. (5) around their steady state. The trivial steady state corresponds to the disordered state with $m^{\mu}_{s}=0$, $q^{ab}_{s}=0$, and $q^{\mu\mu}_{s}\equiv q_{s}(J,K)\neq 0$ which is explicitly given in [55]. Up to first order, small perturbations $\delta\mathbf{m}(t)\equiv(\delta m^{a}(t),\delta m^{b}(t))$ decouple from perturbations $\delta\mathbf{q}(t)\equiv(\delta q^{aa}(t),\delta q^{bb}(t),\delta q^{ab}(t))$, and we obtain the linear equation $\tau d\delta\mathbf{m}(t)/dt=\mathbf{M}(q_{s};J,K)\delta\mathbf{m}(t)$ (see [55] for the linear equation for $\delta\mathbf{q}(t)$). The elements of the linear stability matrix, $M_{\mu\nu}(q_{s};J,K)$, read
$\!\!M_{aa}$ $=M_{bb}=\frac{q_{s}/q_{\rm crit}-1}{1+q_{s}}[1-2\sum\nolimits_{l,n}\!\!\overline{\mathcal{P}}^{+}_{l}\tanh(U^{a}_{l,n})],$
$\!\!M_{ab}$ $={-}M_{ba}{=}\sum\nolimits_{l,n}\!(2n{-}1)[\overline{\mathcal{P}}^{+}_{l}{+}\overline{\mathcal{P}}^{-}_{l}]\tanh(U^{a}_{l,n}),$ (7)
where $\overline{\mathcal{P}}^{\pm}_{l}(q_{s})\equiv\mathcal{C}^{z}_{l}(1\mp q_{s})^{z-l}(1\pm q_{s})^{l}$ are the probabilities (6) evaluated at steady state values, and we introduced the critical local order $q_{\rm crit}\equiv 1/(z-1)$, which depends on the coordination number of the lattice. The solution of the linear stability equation can be written as $\mathbf{\delta m}(t)=\sum_{k=\pm}\mathcal{A}_{k}\mathrm{e}^{\lambda_{k}t/\tau}\boldsymbol{\nu}_{k}$, where $\mathcal{A}_{\pm}$ are set by the initial conditions, $\boldsymbol{\nu}_{\pm}$ are the eigenvectors of $\mathbf{M}(q_{s};J,K)$ and $\lambda_{\pm}(q_{s};J,K)$ the corresponding eigenvalues
$\lambda_{\pm}(q_{s};J,K)=M_{aa}(q_{s};J,K)\pm{\rm i}M_{ab}(q_{s};J,K),$ (8)
$\rm i$ being the imaginary unit. Since ${M_{ab}(q_{s};J,K\neq 0)\neq 0}$ (see proof in [55]), the eigenvalues are complex for $K\neq 0$ (see dashed lines in Fig. 1b), resulting in oscillatory perturbations. The Hopf bifurcation [61], also called type-${\rm II}_{\rm o}$ instability [62], occurs when complex conjugate eigenvalues transit the imaginary axis in the complex plane. According to Eq. (8) this occurs when $M_{aa}(q_{s};J,K)=0$ implying ${q_{s}(J,K)=q_{\rm crit}}$ as seen from Eq. (7). The Hopf bifurcation is thus set by the critical value $q_{\rm crit}$ for local order, and for $q_{s}>q_{\rm crit}$ we have ${\rm Re}(\lambda_{\pm})>0$ (solid line in Fig. 1b). In other words, when spins on the respective lattices are sufficiently aligned, a transition to an oscillatory state occurs (Fig. 1c,d). The critical value of the local order that determines the onset of coherent oscillations is our second main result that generalizes to other approximation schemes beyond the mean field approximation (see Appendix B). After sufficient local order is attained within the lattices, the frustration due to the nonreciprocal coupling gives rise to coherent oscillations: for $K>0$ a spin $\sigma^{a}_{i}$ wants to align with $\sigma^{b}_{i}$ that in turn tends to misalign with $\sigma^{a}_{i}$. This dynamical frustration results in an oscillatory motion of the order parameters [15]. Notably, for the one-dimensional lattice ($z=2$) we find $q_{\rm crit}=1$, which, in contrast to the mean-field prediction [48] (see also Appendix A), correctly implies the non-existence of a Hopf bifurcation.
Figure 2: (a,b) Bifurcation diagram of the magnetization $(m^{a}_{s},m^{b}_{s})$ in the absence (a; $K{=}0$) and presence (b; $K{=}0.01$) of nonreciprocal coupling as a function of $J$. Inset of (b): Magnification of the bifurcation diagram around the Hopf bifurcation projected onto $(m^{b}_{s},J)$-plane (see black box). (c) Temporal evolution of the magnetization close to the degenerate saddle-node-infinite-period (SNIPER) bifurcation at $(J,K){=}(0.35991,0.01)$. The magnetization oscillates coherently between $4$ ghost states ${\rm A}{\rightarrow}{\rm B}{\rightarrow}{\rm C}{\rightarrow}{\rm D}$ (for $K{<}0$ the direction is reversed), which eventually terminate in the $4$ respective stable branches at and beyond the SNIPER bifurcation. The life-time $\mathcal{T}$ in these ghost states diverges upon approaching the SNIPER bifurcation at $J_{\rm SNP}(K)$ from the left as $\lim_{J\uparrow J_{\rm SNP}(K)}\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2}$. (d) Phase diagram of $m^{\mu}_{s}$: the gray region depicts the regime of coherent oscillations, with the black lines indicating isolines with fixed oscillation period, and the blue line indicating the Hopf bifurcations where $q_{s}{=}q_{\rm crit}{\equiv}1/(z{-}1)$; in the light green region the magnetization is stationary and nonzero. In all panels we consider $z=4$.
Non-Linear analysis.—Going beyond linear stability, we perform a non-linear analysis of Eqs. (5) via numerical continuation [63]. The resulting bifurcation diagrams are shown in Fig. 2(a,b) and the complete phase diagram in Fig. 2(d), which we now explain in detail.
We start with the non-interacting case with $K=0$ (see Fig. 2a). For small values of $J$, there exists only one (trivial) stable steady state with $m^{\mu}_{s}=0$, as explained in the previous section. Increasing $J$ to $\ln{(z/(z-2))}/2$ we find a pitchfork bifurcation (red dot in Fig. 2a), which coincides with $q_{s}(J,0)=q_{\rm crit}$, and beyond which the trivial stable state becomes unstable. At the pitchfork bifurcation, $4$ stable branches (blue lines in Fig. 2a) and $4$ unstable branches (red dashed lines in Fig. 2a) emerge. Unstable branches have zero magnetization in one of the lattices, while stable branches exhibit nonzero equilibrium magnetization due to spontaneously broken symmetry in both lattices.
Upon setting $K\neq 0$, the pitchfork bifurcation turns into a Hopf bifurcation (blue dot in Fig. 2b), whose $J$-value depends on $K$ through the relation $q_{s}(J,K)=q_{\rm crit}$ (blue line in Fig. 2d; see [55] for explicit results). Increasing $J$ beyond the Hopf bifurcation, there is a regime with coherent oscillations (gray area in Fig. 2b,d). In Fig. 2d we identify the isolines of fixed period of the oscillations (black lines). Upon further increasing $J$ at fixed $K$, we observe a non-linear transition from coherent oscillations to a non-zero stationary magnetization (light green region in Fig. 2d), which is set by a degenerate saddle-node-infinite-period (SNIPER) bifurcation (green dots in Fig. 2b). Approaching the SNIPER bifurcation from below, the magnetization in both lattices start to oscillate between four long-lived ghost states (see Fig. 2c). These $4$ long-lived states correspond to the virtual configuration, in which the respective lattices exert a “static” magnetic field on each other, and are characterized by a critical slowing down of the dynamics in the vicinity of the impending SNIPER bifurcation. The residence time within these ghost states diverges as (see [55])
$\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2},$ (9)
where $J_{\rm SNP}(K)$ is the $J$-value of the SNIPER bifurcation at a given $K$ (green line in Fig. 2d). At the SNIPER bifurcation, one unstable and stable branch emerge, resulting in $4$ unstable branches (red dashed lines in Fig. 2b) and $4$ stable branches (blue lines in Fig. 2b). Note that these stable states with non-zero stationary magnetization exist also in the finite square lattice system (see proof in [55]).
The bifurcation diagram obtained with the mean-field approximation has similar qualitative features as in Fig. 2a,b (i.e, a Hopf and SNIPER bifurcation), however, the phase diagram displays a constant Hopf line at a fixed $J$, independent of $K$ (see [48] and Fig. A1). In Fig. 2d we see that the Hopf line with the BG approximation is $K$-dependent, closely resembling the empirical phase diagram on the qubic lattice (see [48]).
Spatio-temporal dynamics.—The results in Fig. 1d-f reveal a high degree of local order in the coherent oscillatory regime. To systematically analyze the spatio-temporal patterns in states with coherent oscillations, we performed discrete-time Monte-Carlo simulations of the nonreciprocal Ising system with $2\times N\approx 3\times 10^{3}$ spins on the all-to-all ($z{=}N$), the $z{=}4$ Bethe, and the square ($z{=}4$) lattices with periodic boundary conditions. The respective systems are described in detailed in [55]. For a consistent notion of “spatial scale” on all lattices, we perform a graph-spectral analysis [64].
Figure 3: (a-c) Spectral density $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ of modes of coherent oscillations (Eq. (10)) on the (a) mean-field, (b) Bethe, and (c) square lattice (see [55] for videos of the dynamics on the lattices). Note that the nonzero eigenvalues of the mean-field lattice are degenerate, which explains the white region in (a). The values for $J$ are chosen in the respective oscillatory regimes (see [55] for details), and $K=0.3$ for all three lattices. $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ is averaged over $500$ independent trajectories, and $\omega_{0}={\rm argmax}\langle|\hat{\Psi}^{a}_{1}(\omega)|\rangle$ is the natural oscillation frequency. Resonances are visible at multiplies of $2\omega_{0}$, since we ignore the sign of projections. Insets: Temporal development of the projected microscopic states onto the eigenvectors of the Laplacian matrix of respective lattices. Colors denote three selected eigenvalues (and thus spatial scales) indicated by the dots in the main plot.
Let $\mathbf{L}$ be the $N{\times}N$ symmetric Laplacian matrix of one of the above graphs with elements $L_{ii}{=}z$, and $L_{ij}{=}-1$ when spin $i$ and $j$ are connected and $L_{ij}{=}0$ otherwise. The Laplacian has $N$ orthonormal eigenvectors $\mathbf{L}\boldsymbol{\psi}_{k}{=}l_{k}\boldsymbol{\psi}_{k},\ k\in\{1,...,N\}$ with corresponding eigenvalues $l_{k}$ ordered as $l_{1}{\leq}l_{2}{\leq}\ldots{\leq}l_{N}$. The lowest eigenvalue, corresponding to $\boldsymbol{\psi}_{1}{=}N^{-1/2}(1,\ldots,1)^{\rm T}$, vanishes, i.e., $l_{1}{=}0$ [65]. For the mean-field lattice, all remaining $N{-}1$ eigenvalues are degenerate, $l_{2}=\ldots=l_{N}=N$ [66], but not for the square and Bethe lattices (see [55]).
We express the microscopic state of the lattice $\mu$ at time-step $n\in\{0,...,n_{\rm max}\}$ as a column vector $\boldsymbol{\sigma}^{\mu}(n)=(\sigma^{\mu}_{1}(n),...,\sigma^{\mu}_{N}(n))^{T}$ and project it onto the respective eigenvectors, $\Psi^{\mu}_{k}(n)\equiv\boldsymbol{\psi}_{k}^{T}\boldsymbol{\sigma}^{\mu}(n)$. These spatial modes are shown as insets in Fig. 3a-c, where one can already see that for the Bethe and square lattice oscillations are pronounced on large scales (i.e. for small $k$) and severely suppressed on small scales (large $k$). In the mean-field system, the magnitude of (stochastic) oscillations is equal on all scales (all $k$) since the spectrum of $\mathbf{L}$ is trivial.
To unravel the spatio-temporal structure, we compute the spectral density via the discrete Fourier transform,
$\langle|\hat{\Psi}^{\mu}_{k}(\omega)|\rangle\equiv\biggl{\langle}\biggl{|}\sum\nolimits_{n=0}^{n_{\rm max}}|\Psi^{\mu}_{k}(n)|\mathrm{e}^{-{\rm i}2\pi\omega n/n_{\rm max}}\biggr{|}\biggr{\rangle},$ (10)
[…]
Conclusion.—We explained the collective dynamics of the nonreciprocal Ising system on the level of both, local and global order. A critical threshold magnitude of local order within the respective lattices was found to control the emergence of coherent oscillations of the global order parameter. Upon increasing interactions, ghost states emerge and the residence time in either of them eventually diverges, giving rise to a non-equilibrium steady state via a saddle-node-infinite-period bifurcation. Strikingly, during coherent oscillations of global order, a high degree of local order is preserved in nearest-neighbor-interacting systems. This implies non-trivial spatio-temporal correlations between spins, confirmed by a spectral-density maximum at large-scale low-frequency modes. In stark contrast, on the mean-field (all-to-all) lattice there is no distinction between different spatial modes, annihilating any notion of spatial structure. Thus, accounting for nearest-neighbor correlations is essential for a correct understanding of dynamics of nonreciprocal matter with short, or more generally finite, range of interactions.
Our work provides a comprehensive microscopic understanding of dynamic collective phenomena in nonreciprocal matter without conservation laws relevant, e.g. for enzyme clustering [67, 68], where the nonreciprocal coupling relates to an attractive $K>0$ or repulsive $K<0$ interaction between two enzymes where the product of one is the substrate of the other enzyme. The magnetizations $m^{\mu}(t)$ in turn reflect time-dependent concentrations of product and substrates via $\phi^{\mu}(t)=(1+m^{\mu}(t))/2\in[0,1]$.
What remains elusive are multiple ($>2$) coupled lattices, spatially heterogeneous/extended systems [69], as well as the thermodynamic cost of dynamical states and bifurcations [53, 70, 71, 72]. Moreover, considering the relevance of conservation laws [73, 2, 41], it will be essential to develop a theoretical framework for the nonreciprocal Ising model with Kawasaki dynamics. These will be addressed in future works.
Acknowledgments.— Financial support from the German Research Foundation (DFG) through the Heisenberg Program Grants GO 2762/4-1 and GO 2762/5-1 (to AG) is gratefully acknowledged.
Appendix A: Mean-field approximation.—A less accurate technique to obtain approximate evolution equations is the mean-field (MF) approximation [48], where one makes the rudimentary (uncontrolled) assumption
$\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle\approx\tanh{\langle\Delta E^{\mu}_{i}/2\rangle},$ (A1)
yielding the evolution equations
$\tau\frac{dm^{a}(t)}{dt}+m^{a}(t)$ $=\tanh{(zJ_{a}m^{a}(t)+K_{a}m^{b}(t))},$
$\tau\frac{dm^{b}(t)}{dt}+m^{b}(t)$ $=\tanh{(zJ_{b}m^{b}(t)+K_{b}m^{a}(t))},$ (A2)
which are exact on the fully connected mean-field lattice, where the local order is trivial (i.e. there is no sense of “local”), $q^{\mu\nu}(t)=m^{\mu}(t)m^{\nu}(t)$, and therefore $\mathcal{C}^{\mu\nu}(t)=0$. A linear stability analysis around the trivial steady-state $m^{\mu}_{s}=0$ for $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$ leads to the linear stability equation $\tau d\delta\mathbf{m}(t)/dt=\mathbf{M}^{\rm MF}(J,K)\delta\mathbf{m}(t)$ with elements
$M^{\rm MF}_{aa}$ $=M^{\rm MF}_{bb}=zJ-1,$ (A3)
$M^{\rm MF}_{ab}$ $=-M^{\rm MF}_{ba}=K.$ (A4)
The eigenvalues of $\mathbf{M}^{\rm MF}$ are given by
$\lambda^{\rm MF}_{\pm}(J,K)=(zJ-1)\pm iK.$ (A5)
Hence, the Hopf bifurcation occurs at $J=1/z$ and $K\neq 0$, such that ${\rm Re}(\lambda^{\rm MF}_{\pm})=0$ and ${\rm Im}(\lambda^{\rm MF}_{\pm})\neq 0$. This corresponds to a straight vertical line in the $(J,K)$-plane, as shown in Fig. A1a, which is unphysical for short range interactions. Notably, in the MF approximation we do not observe a critical value for local order, which is present in the more accurate BG approximation.
Figure A1: Phase diagram for global order $m^{\mu}_{s}$ obtained with the mean-field (MF) approximation (a) and the monomer approximation (b) for the perfect nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$. In the MF approximation the Hopf bifurcation (blue line) is set by $J=1/z$. In the monomer approximation, the Hopf bifurcations (blue line) is set by a critical local order $q_{s}=1/z$, which is more similar to the Bethe-Guggenheim approximation where $q_{s}=1/(z-1)$ (see Fig. 2d).
Table 1: Overview of approximation techniques and their dynamical equations
approx. technique $m^{\mu}(t)$ $q^{\mu\nu}(t)$ Hopf
MF (A2) $=m^{\mu}(t)m^{\nu}(t)$ $J{=}1/z$
monomer (5)+(B2) (5)+(B2); slaved by $m^{\mu}(t)$ $q_{s}{=}1/z$
BG (5)+(6) (5)+(6); not slaved $q_{s}{=}1/(z{-}1)$
Appendix B: Monomer approximation.—Another approximation technique we developed in this work is what we call the “monomer approximation”. It is more accurate than the MF yet less accurate than the BG approximation.
The conceptual difference between the MF on the one hand, and the monomer and BG approximations on the other hand lies in the treatment of the average $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle$. Whereas the MF approximation simply moves the average to the argument as shown in Eq. (A1), the monomer and BG approximations use the fact that the value of $\Delta E^{\mu}_{i}/2$ lies in an enumerable set given by $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$ with $l\in\{0,..,z\}$ and $n\in\{0,1\}$. This allows for an explicit summation
$\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (B1)
where only the probability $\mathcal{P}^{\mu}_{l,n}(t)$ has to approximated.
Similarly to the BG approximation, the resulting evolution equations in the monomer approximation are governed by Eqs. (5), but the time-dependent probabilities are different and read (the derivation is given in [55])
$\mathcal{P}^{a}_{l,n}(t)=\frac{2\mathcal{C}^{z}_{l}(1+m^{a}(t))^{l}(1+m^{b}(t))^{n}}{(1-m^{a}(t))^{l-z}(1-m^{b}(t))^{n-1}},$ (B2)
and $\mathcal{P}^{b}_{l,n}(t)$ is obtained by replacing $m^{a}(t)$ with $m^{b}(t)$ in Eq. (B2). Since $\mathcal{P}^{a}_{l,n}(t)$ is independent of the local order $q^{\mu\nu}(t)$, this implies that $q^{\mu\nu}(t)$ is slaved by $m^{\mu}(t)$, however, $q^{\mu\nu}(t)\neq m^{\mu}(t)m^{\nu}(t)$. Performing a linear stability analysis around the trivial steady-state $m^{\mu}_{s}=0$ for $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, we obtain the linear stability equation $\tau d\delta\mathbf{m}(t)/dt=\mathbf{\hat{M}}(J,K)\delta\mathbf{m}(t)$ where the elements of $\mathbf{\hat{M}}(J,K)$ are given by
$\hat{M}_{aa}$ $=\hat{M}_{bb}=2\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)\mathcal{C}^{z}_{l}\tanh{(U^{a}_{l,n})}-1,$ (B3)
$\hat{M}_{ab}$ $=-\hat{M}_{ba}=2\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)\mathcal{C}^{z}_{l}\tanh{(U^{a}_{l,n})}.$ (B4)
Interestingly, Eq. (B3) can be expressed in terms of the steady-state value of the local order in the monomer approximation, i.e., $\hat{M}_{aa}(J,K)=zq_{s}(J,K)-1$. Therefore, the eigenvalues of $\mathbf{\hat{M}}(J,K)$ are neatly given by
$\hat{\lambda}_{\pm}(J,K)=(zq_{s}(J,K)-1)\pm i\hat{M}_{ab}(J,K).$ (B5)
Herefrom it follows that the Hopf bifurcation occurs at $q_{s}=1/z$ and $K\neq 0$, such that ${\rm Re}(\hat{\lambda}_{\pm})=0$ and ${\rm Im}(\hat{\lambda}_{\pm})\neq 0$. Hence, as within the BG approximation, we also find the existence of a critical local order in the monomer approximation. Contrary to the MF approximation, the Hopf line is not a straight vertical line in the $(J,K)$-plane, as shown in Fig. A1b.
Appendix C: Local and global order in various approximations.—To provide a concise overview of the various approximation techniques, we summarize in Table 1 the respective evolution and conditions for the Hopf bifurcation.
Supplementary Material for: Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law
Kristian Blom1, Uwe Thiele2,3,4, and Aljaž Godec1
1Mathematical bioPhysics Group, Max Planck Institute for Multidisciplinary Sciences, 37077 Göttingen, Germany
2Institute of Theoretical Physics, University of Münster, Münster 48149, Germany
3Center for Nonlinear Science (CeNoS), University of Münster, Münster 48149, Germany
4Center for Multiscale Theory and Computation (CMTC), University of Münster, Münster 48149, Germany
In this Supplemental Material, we provide derivations and further details about the results shown in the Letter. Furthermore, in Sec. S6 we prove that static global order in the square-lattice Ising model persists at least under small perturbations in the nonreciprocal coupling $0<|K|\leq\mathcal{O}(1/N)$ for any finite but arbitrarily large system size $N\in[m,\infty)$ for some sufficiently large $m$.
Appendix S1 Setup: Dual lattice topology and interactions
Consider a pair of lattices, denoted by $\mu=a,b$, each having a coordination number $z$ and periodic boundary conditions. On each lattice there are $N$ spins that can be in two states $\sigma^{\mu}_{i}=\pm 1$, with $i\in\{1,...,N\}$ enumerating the spin’s location. Each spin interacts with its $z$ nearest-neighbors on the same lattice and one spin at the equivalent position on the opposing lattice. The local interaction energy of spin $i$ on lattice $\mu$ can be written as (see also [1])
$E^{\mu}_{i}=-J_{\mu}\sigma^{\mu}_{i}\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}-K_{\mu}\sigma^{a}_{i}\sigma^{b}_{i},$ (S1)
where $\langle i|j\rangle$ denotes a sum over nearest-neighbors $j$ of spin $i$. Throughout, we will express energies in units of $k_{\rm B}T$. The parameters $J_{\mu}$ and $K_{\mu}$ denote the coupling within one and between the two lattices, respectively. When $K_{a}\neq K_{b}$, equivalent spins in the opposing lattices interact nonreciprocally.
Appendix S2 Glauber dynamics
We now specify the dynamics. We consider single spin-flip dynamics according to Glauber [2]. Let $P(\boldsymbol{\sigma};t)$ be the probability at time $t$ to be in state $\boldsymbol{\sigma}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$. This state can be reached by a single spin flip from the $2N$ distinct states $\boldsymbol{\sigma}^{\prime}_{\mu,i}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,-\sigma^{\mu}_{i},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$, which enters the master equation as follows
$\frac{{\rm d}P(\boldsymbol{\sigma};t)}{{\rm d}t}=\sum_{\mu=a,b}\sum_{i=1}^{N}[w^{\mu}_{i}(-\sigma^{\mu}_{i})P(\boldsymbol{\sigma}^{\prime}_{\mu,i};t)-w^{\mu}_{i}(\sigma^{\mu}_{i})P(\boldsymbol{\sigma};t)].$ (S2)
The respective transition rates $w^{\mu}_{i}(\sigma^{\mu}_{i})$ to flip a single spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$ can be specified uniquely by limiting the interactions to nearest neighbors, imposing isotropy in position space, and requiring that for $K_{a}=K_{b}$ the transition rates obey detailed-balance. These physical restrictions then lead to the general result
$w^{\mu}_{i}(\sigma^{\mu}_{i})=(1/2\tau)\left[1-\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\right],$ (S3)
where $\Delta E^{\mu}_{i}$ is the change in energy on the $\mu=a\lor b$ lattice after flipping spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau$ is an intrinsic timescale to attempt a single spin-flip. In a discrete-time setting the intrinsic flip rate $1/\tau$ is replaced by the probability to flip a single spin, which is $1/2N$ (see also Fig. S2). According to Eq. (S1), the change in energy after such a spin-flip is given by
$\Delta E^{\mu}_{i}=-2E^{\mu}_{i}.$ (S4)
From the master equation (S2), we can directly derive equations for the first two moments of single-spin values (see also Eqs. (28) and (29) in [2])
$\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\rangle}{{\rm d}t}+\langle\sigma^{\mu}_{i}(t)\rangle$ $=\langle\sigma^{\mu}_{i}(t)\tanh{(\Delta E^{\mu}_{i}/2)}\rangle,$ (S5)
$\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)\rangle}{{\rm d}t}+2\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)\rangle$ $=\langle\sigma^{\mu}_{i}(t)\sigma^{\nu}_{j}(t)[\tanh{(\Delta E^{\mu}_{i}/2)}+\tanh{(\Delta E^{\nu}_{j}/2)}]\rangle,$ (S6)
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Eqs. (S5)-(S6) are exact (but not yet closed) equations for the first two moments evolving under Glauber dynamics, and will serve as our starting point to derive equations for the global and local order parameters.
Appendix S3 Differential equations for the global and local order
Recall from the Letter that the global and local order are defined as
$m^{\mu}(t)$ $\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{\mu}_{i}(t)\rangle,$ (S7)
$q^{\mu\mu}(t)$ $\equiv\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma_{i}^{\mu}(t)\sigma_{j}^{\mu}(t)\rangle,$ (S8)
$q^{ab}(t)$ $\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma_{i}^{a}(t)\sigma_{i}^{b}(t)\rangle.$ (S9)
Here, we rewrite Eqs. (S5)-(S6) in terms of differential equations for (S7)-(S9), which we do in two steps:
  1. 1.
    First, we sum Eqs. (S5)-(S6) over all spins and spin pairs. Upon doing this, the l.h.s. of Eqs. (S5)-(S6) directly transforms into
    $\frac{1}{N}\sum_{i=1}^{N}\left(\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\rangle}{{\rm d}t}+\langle\sigma^{\mu}_{i}(t)\rangle\right)\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =\tau\frac{{\rm d}m^{\mu}(t)}{{\rm d}t}+m^{\mu}(t),$ (S10)
    $\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\left(\tau\frac{{\rm d}\langle\sigma^{\mu}_{i}(t)\sigma^{\mu}_{j}(t)\rangle}{{\rm d}t}+2\langle\sigma^{\mu}_{i}(t)\sigma^{\mu}_{j}(t)\rangle\right)=\tau\frac{{\rm d}q^{\mu\mu}(t)}{{\rm d}t}+2q^{\mu\mu}(t),$ (S11)
    $\frac{1}{N}\sum_{i=1}^{N}\left(\tau\frac{{\rm d}\langle\sigma^{a}_{i}(t)\sigma^{b}_{i}(t)\rangle}{{\rm d}t}+2\langle\sigma^{a}_{i}(t)\sigma^{b}_{i}(t)\rangle\right)\ \ \ \ \ \ \ \ =\tau\frac{{\rm d}q^{ab}(t)}{{\rm d}t}+2q^{ab}(t).$ (S12)
  2. 2.
    Next, we need to evaluate the r.h.s. of Eqs. (S5)-(S6) after summation over all spins and spin pairs. To do this, we make use of the fact that $\Delta E^{\mu}_{i}$ can take on a discrete (enumerable) set of values. Let us consider a spin with $l\in\{0,1,...,z\}$ neighboring up spins on the same lattice, and $n\in\{0,1\}$ neighboring up spins on the opposing lattice. We want to compute the change in energy upon flipping this spin. Based on Eq. (S4) we can parameterize the change in energy after flipping a spin with such an environment as follows,
    $\Delta E^{\mu}_{i}=2\sigma^{\mu}_{i}U^{\mu}_{l,n},$ (S13)
    where
    $U^{\mu}_{l,n}=[2l-z]J_{\mu}+[2n-1]K_{\mu}.$ (S14)
    This gives a direct parameterization of the energy change in terms of the number of neighboring up spins on the same $(l)$ and opposing $(n)$ lattice. Using this parameterization, we can evaluate the r.h.s. of Eqs. (S5)-(S6) using the following identities
    $\!\!\!\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{\mu}_{i}\tanh{(\Delta E^{\mu}_{i}/2)}\rangle$ $=\frac{1}{N}\sum_{i=1}^{N}\langle\tanh{(U^{\mu}_{l,n})}\rangle=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (S15)
    $\!\!\!\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma^{\mu}_{i}\sigma^{\mu}_{j}\tanh{(\Delta E^{\mu}_{i}/2)}\rangle$ $=\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma^{\mu}_{j}\tanh{(U^{\mu}_{l,n})}\rangle=\frac{2}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (S16)
    $\!\!\!\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{a}_{i}\sigma^{b}_{i}\tanh{(\Delta E^{a}_{i}/2)}\rangle$ $=\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{b}_{i}\tanh{(U^{a}_{l,n})}\rangle=\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{a}_{l,n})}.$ (S17)
For the first equality in Eqs. (S15)-(S17) we used that $\tanh{(\Delta E^{\mu}_{i}/2)}=\tanh{(\sigma^{\mu}_{i}U^{\mu}_{l,n})}=\sigma^{\mu}_{i}\tanh{(U^{\mu}_{l,n})}$, together with $(\sigma^{\mu}_{i})^{2}=1$. For the second equality, we used that terms like $\langle\tanh{(U^{\mu}_{l,n})}\rangle$ represent a weighted sum over all possible combinations of the possible values that $\tanh{(U^{\mu}_{l,n})}$ can attain. The weights are given by the time-dependent probability $\mathcal{P}^{\mu}_{l,n}(t)$ to find an up or down spin with a specific local environment. By definition, this probability is normalized as,
$\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)=1.$ (S18)
Combining Eqs. (S10)-(S12) and (S15)-(S17) we obtain
$\tau\frac{dm^{\mu}(t)}{dt}+m^{\mu}(t)$ $=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$ (S19)
$\tau\frac{dq^{\mu\mu}(t)}{dt}+2q^{\mu\mu}(t)$ $=\frac{2}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$ (S20)
$\tau\frac{dq^{ab}(t)}{dt}+2q^{ab}(t)$ $=\sum_{\mu=a,b}\sum_{l=0}^{z}\sum_{n=0}^{1}(2n{-}1)\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$ (S21)
corresponding to Eq. (5) in the Letter. Our next aim is to obtain closed-form expressions for $\mathcal{P}^{\mu}_{l,n}(t)$, for which we consider two approximation techniques: the “monomer” approximation, and the Bethe-Guggenheim (BG) approximation.
S3.1 The “monomer” approximation
We start with the lowest-order approximation, where we neglect nearest-neighbor correlations in the probability $\mathcal{P}^{\mu}_{l,n}$. Note, however, that already this level of approximation goes beyond the mean-field approach in [1]. Suppose we want to know the probability of finding an up or down spin on the $a$ lattice which has $l$ nearest up neighbors on the $a$ lattice and $n$ nearest up neighbors on the $b$ lattice. On the monomer level we assume perfect mixing between the up and down spins, resulting in
$\mathcal{P}^{a}_{l,n}=\underbrace{\left[\binom{N^{a}_{+}}{l}\binom{N^{a}_{-}}{z-l}/\binom{N^{a}_{+}+N^{a}_{-}}{z}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}l\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{b}_{+}}{n}\binom{N^{b}_{-}}{1-n}/\binom{N^{b}_{+}+N^{b}_{-}}{1}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}n\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ b\ {\rm lattice}\end{subarray}\end{subarray}},$ (S22)
where $N^{\mu}_{+}$ is the total number of up spins on the $\mu$ lattice, and similarly $N^{\mu}_{-}$ the total number of down spins. The same reasoning applies to the probability $\mathcal{P}^{b}_{l,n}$. Note that $N^{\mu}_{\pm}$ is generally time-dependent, but for simplicity we omit the subscript $t$. To relate $N^{\mu}_{\pm}$ to the global order we use the following relations (omitting the time dependence subscript $t$)
$m^{\mu}$ $=(N^{\mu}_{+}-N^{\mu}_{-})/N,$ (S23)
$N$ $=N^{\mu}_{+}+N^{\mu}_{-},$ (S24)
from which follows that
$N^{\mu}_{\pm}$ $=N(1\pm m^{\mu})/2.$ (S25)
Inserting this back into Eq. (S22) and taking the thermodynamic limit, i.e. the scaling limit $N\rightarrow\infty$ while keeping $\mathbf{m}(t)\equiv(m^{a}(t),m^{b}(t))$ fixed, we can make use of the following result for the binomial coefficients
$\lim\limits^{m^{\mu}={\rm const.}}_{N\rightarrow\infty}\binom{N(1\pm m^{\mu})/2}{l}\simeq\frac{(N(1\pm m^{\mu})/2)^{l}}{l!},\ {\rm for}\ l\in\mathbb{N}$ (S26)
where $\simeq$ stands for asymptotic equality. Inserting Eq. (S26) into Eq. (S22), and restoring the explicit time-dependence, we finally obtain (note that the $N$ dependence cancels out)
$\mathcal{P}^{a}_{l,n}(t)$ $\simeq\frac{2\mathcal{C}^{z}_{l}(1+m^{a}(t))^{l}(1+m^{b}(t))^{n}}{(1-m^{a}(t))^{l-z}(1-m^{b}(t))^{n-1}},$ (S27)
$\mathcal{P}^{b}_{l,n}(t)$ $\simeq\frac{2\mathcal{C}^{z}_{l}(1+m^{b}(t))^{l}(1+m^{a}(t))^{n}}{(1-m^{b}(t))^{l-z}(1-m^{a}(t))^{n-1}},$ (S28)
where
$\mathcal{C}^{z}_{l}\equiv\frac{1}{2^{z+2}}\binom{z}{l}.$ (S29)
Hence, in the monomer approximation, we find that the probability $\mathcal{P}^{\mu}_{l,n}(t)$ only explicitly depends on the global order $\mathbf{m}(t)$, and not on the local order $\mathbf{q}(t)\equiv(q^{aa}(t),q^{bb}(t),q^{ab}(t))$. Upon inserting Eq. (S27) into Eqs. (S19)-(S21) we obtain a closed set of five non-linear differential equations, where, notably, the dynamical equations for $\mathbf{q}(t)$ are slaved by the solutions for $\mathbf{m}(t)$, as the r.h.s. of Eqs. (S19)-(S21) do not depend on $\mathbf{q}(t)$.
S3.1.1 Difference with respect to mean-field approximation
Finally, let us point out the crucial difference between the mean-field on the one hand and the monomer and BG approximation on the other hand, which lies in the treatment of the averaging of the term $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle$. Whereas the mean-field approximation (uncontrollably) moves the average to the argument, i.e., $\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle\approx\tanh{\left(\langle\Delta E^{\mu}_{i}/2\rangle\right)}$, the monomer and BG approximations use the fact that the value of $\Delta E^{\mu}_{i}/2$ lies in an enumerable set given by $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$ with $l\in\{0,..,z\}$ and $n\in\{0,1\}$. This allows for an explicit summation
$\langle\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\rangle=\sum_{l=0}^{z}\sum_{n=0}^{1}\mathcal{P}^{\mu}_{l,n}(t)\tanh{(U^{\mu}_{l,n})},$ (S30)
where only the probability $\mathcal{P}^{\mu}_{l,n}(t)$ is approximated according to Eq. (S27)-(S28) in the monomer approximation.
S3.2 The Bethe-Guggenheim “pair” approximation
In the previous method, we did not take into account nearest-neighbor correlations in the evaluation of $\mathcal{P}^{\mu}_{l,n}(t)$, which we aim to include in this Section. Let us focus on the probability of picking a spin on the $a$ lattice with a given specific local environment. The same reasoning will also apply for picking a spin on the $b$ lattice. First, we split the probability in a contribution of picking an up or down spin
$\mathcal{P}^{a}_{l,n}(t)=\mathcal{P}^{a+}_{l,n}(t)+\mathcal{P}^{a-}_{l,n}(t),$ (S31)
where, e.g., $\mathcal{P}^{a+}_{l,n}(t)$ is the probability of picking an up spin with $l$ neighboring up spins on the $a$ lattice and $n$ neighboring up spins on the $b$ lattice at time $t$. On the BG level we assume ideal mixing of nearest-neighbor spin pairs, resulting in the following expression
$\mathcal{P}^{a+}_{l,n}$ $=\underbrace{[N^{a}_{+}/N]}_{\begin{subarray}{c}\rm probability\\ \begin{subarray}{c}\rm for\ up\ spin\\ {\rm on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{aa}_{++}}{l}\binom{N^{aa}_{+-}/2}{z-l}/\binom{N^{aa}_{++}+N^{aa}_{+-}/2}{z}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}l\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{ab}_{++}}{n}\binom{N^{ab}_{+-}}{1-n}/\binom{N^{ab}_{++}+N^{ab}_{+-}}{1}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}n\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ b\ {\rm lattice}\end{subarray}\end{subarray}},$ (S32)
$\mathcal{P}^{a-}_{l,n}$ $=\underbrace{[N^{a}_{-}/N]}_{\begin{subarray}{c}\rm probability\\ \begin{subarray}{c}\rm for\ down\ spin\\ {\rm on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{aa}_{-+}/2}{l}\binom{N^{aa}_{--}}{z-l}/\binom{N^{aa}_{-+}/2+N^{aa}_{--}}{z}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}l\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ a\ {\rm lattice}\end{subarray}\end{subarray}}\times\underbrace{\left[\binom{N^{ab}_{-+}}{n}\binom{N^{ab}_{--}}{1-n}/\binom{N^{ab}_{-+}+N^{ab}_{--}}{1}\right]}_{\begin{subarray}{c}\rm probability\ for\\ \begin{subarray}{c}n\ {\rm\ neighboring\ up}\\ {\rm spins\ on\ the}\ b\ {\rm lattice}\end{subarray}\end{subarray}},$ (S33)
where, for example, $N^{ab}_{+-}$ is the total number of nearest-neighbor spin pairs with an up spin on the $a$ lattice and a down spin on the $b$ lattice. To relate $N^{ab}_{+-}$ and the other spin pair numbers to the global and local order, we make use of the following exact relations for periodic lattices
$2N^{\mu\mu}_{\pm\pm}+N^{\mu\mu}_{+-}$ $=zN^{\mu}_{\pm},$ (S34)
$N^{ab}_{\pm\pm}+N^{ab}_{\pm\mp}$ $=N^{a}_{\pm},$ (S35)
in combination with Eq. (S25). Furthermore, we use the definition of local order (see Eqs. (S8)-(S9)) to write
$q^{\mu\mu}$ $=2(N^{\mu\mu}_{++}+N^{\mu\mu}_{--}-N^{\mu\mu}_{+-})/zN=1-4N^{\mu\mu}_{+-}/zN,$ (S36)
$q^{ab}$ $=(N^{ab}_{++}+N^{ab}_{--}-N^{ab}_{+-}-N^{ab}_{-+})/N$
$=1-2N^{ab}_{+-}/N-2N^{ab}_{-+}/N$
$=1+m^{a}-m^{b}-4N^{ab}_{+-}/N,$ (S37)
where in the last line we used the exact relation
$m^{b}-m^{a}=2(N^{ab}_{-+}-N^{ab}_{+-})/N.$ (S38)
Using the relations (S34)-(S37) we obtain the following result for the spin pairs within the same lattice
$N^{\mu\mu}_{\pm\pm}$ $=(z/8)N(1\pm 2m^{\mu}+q^{\mu\mu}),$ (S39)
$N^{\mu\mu}_{\pm\mp}$ $=(z/4)N(1-q^{\mu\mu}),$ (S40)
and for the spin pairs between the two opposing lattices
$N^{ab}_{\pm\pm}$ $=(1/4)N(1\pm m^{a}\pm m^{b}+q^{ab}),$ (S41)
$N^{ab}_{\pm\mp}$ $=(1/4)N(1\pm m^{a}\mp m^{b}-q^{ab}).$ (S42)
Inserting these relations back into Eqs. (S32)-(S33) and taking the thermodynamic limit $N\rightarrow\infty$ while keeping $\mathbf{m}(t)$ and $\mathbf{q}(t)$ fixed, we obtain (omitting, for convenience, arguments on the l.h.s.)
$\mathcal{P}^{a+}_{l,n}(t)$ $\simeq\frac{\mathcal{C}^{z}_{l}(1+2m^{a}(t)+q^{aa}(t))^{l}(1+m^{a}(t)-m^{b}(t)-q^{ab}(t))^{1-n}}{(1+m^{a}(t))^{z}(1-q^{aa}(t))^{l-z}(1+m^{a}(t)+m^{b}(t)+q^{ab}(t))^{-n}},$ (S43)
$\mathcal{P}^{a-}_{l,n}(t)$ $\simeq\frac{\mathcal{C}^{z}_{l}(1-2m^{a}(t)+q^{aa}(t))^{z-l}(1-m^{a}(t)-m^{b}(t)+q^{ab}(t))^{1-n}}{(1-m^{a}(t))^{z}(1-q^{aa}(t))^{-l}(1-m^{a}(t)+m^{b}(t)-q^{ab}(t))^{-n}},$ (S44)
where $\mathcal{C}^{z}_{l}$ is given in Eq. (S29). Expressions (S43)-(S44) are given in the main Letter in Eq. (6). The expressions for $\mathcal{P}^{b\pm}_{l,n}(t)$ are similar to Eqs. (S43)-(S44) upon interchanging $m^{a}(t)$ and $m^{b}(t)$ and replacing $q^{aa}(t)$ with $q^{bb}(t)$. The difference with the “monomer” approximation (S27) and (S28) is that in the BG approximation the probability also depends explicitly on the local order $\mathbf{q}(t)$. Inserting Eqs. (S43)-(S44) into Eq. (S31), and finally into Eqs. (S19)-(S21), leads to a closure of the five coupled non-linear differential equations.
Appendix S4 Linear stability analysis
Here, we investigate the stability of steady-states of Eqs. (5) in the Letter, in combination with the aforementioned approximation techniques, using linear stability analysis. This allows us to identify the region in parameter space where we have a so-called Hopf bifurcation, which marks the transition from a (non-oscillatory) steady-state to coherent oscillations.
S4.1 The “monomer” approximation
Since in the MF approximation the local order is slaved by the global order, it suffices to consider the linear stability analysis for the global order. Let us consider a small perturbation of the global order around the steady-state value $\textbf{m}_{\rm s}$, i.e., $\mathbf{m}(t)=\mathbf{m}_{\rm s}+\mathbf{\delta m}(t)$. The most interesting steady-state value to consider is given by the disordered state $\mathbf{m}_{\rm s}=0$, in which case the probability (S27)-(S28) can be expanded as
$\mathcal{P}^{a}_{l,n}(t)$ $=2\mathcal{C}^{z}_{l}\left(1+(2l-z)\delta m^{a}(t)+(2n-1)\delta m^{b}(t)\right)+\mathcal{O}(\mathbf{\delta m}^{2}(t)),$ (S45)
$\mathcal{P}^{b}_{l,n}(t)$ $=2\mathcal{C}^{z}_{l}\left(1+(2l-z)\delta m^{b}(t)+(2n-1)\delta m^{a}(t)\right)+\mathcal{O}(\mathbf{\delta m}^{2}(t)).$ (S46)
Inserting this linearized expression back into Eq. (S19), we eventually obtain
$\tau\frac{{\rm d}\mathbf{\delta m}(t)}{{\rm d}t}=\underbrace{\begin{pmatrix}\hat{M}_{aa}&\hat{M}_{ab}\\ \hat{M}_{ba}&\hat{M}_{bb}\end{pmatrix}}_{\mathbf{\hat{M}}(J,K)}\mathbf{\delta m}(t)+\mathcal{O}(\mathbf{\delta m}^{2}(t)),$ (S47)
where the entries of the matrix $\mathbf{\hat{M}}$ read
$\hat{M}_{\mu\mu}(J_{\mu},K_{\mu})$ $=2\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)\mathcal{C}^{z}_{l}\tanh{(U^{\mu}_{l,n})}-1,$ (S48)
$\hat{M}_{ab}(J_{a},K_{a})$ $=2\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)\mathcal{C}^{z}_{l}\tanh{(U^{a}_{l,n})},$ (S49)
$\hat{M}_{ba}(J_{b},K_{b})$ $=2\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)\mathcal{C}^{z}_{l}\tanh{(U^{b}_{l,n})}.$ (S50)
Note that $\hat{M}_{\mu\mu}$ can be written in terms of the steady-state solution for the local order given by Eq. (S20), i.e., $\hat{M}_{\mu\mu}=zq^{\mu\mu}_{s}-1$, which we will make use of in the remaining calculation. Furthermore, the off-diagonals are related to the steady-state of the local order between the lattices, i.e., $\hat{M}_{ab}+\hat{M}_{ba}=2q^{ab}_{s}$. Hence, a perturbation of the global order couples to the steady-state value of the local order. The solution of Eq. (S47) can be written in terms of an eigenmode expansion
$\mathbf{\delta m}(t)=\sum_{k=\pm}\mathcal{A}_{k}\mathrm{e}^{\hat{\lambda}_{k}t/\tau}\boldsymbol{\hat{\nu}}_{k},$ (S51)
where $\mathcal{A}_{\pm}$ are set by the initial conditions, $\hat{\lambda}_{\pm}$ are the eigenvalues of the linear stability matrix
$\hat{\lambda}_{\pm}=\left({\rm tr}(\mathbf{\hat{M}})\pm\sqrt{{\rm tr}(\mathbf{\hat{M}})^{2}-4{\rm det}(\mathbf{\hat{M}})}\right)/2,$ (S52)
with
${\rm tr}(\mathbf{\hat{M}})$ $=z\left(q^{aa}_{s}+q^{bb}_{s}\right)-2,$ (S53)
${\rm det}(\mathbf{\hat{M}})$ $=(zq^{aa}_{s}-1)(zq^{bb}_{s}-1)-\hat{M}_{ab}\hat{M}_{ba},$ (S54)
and the eigenvectors $\boldsymbol{\hat{\nu}}_{\pm}$ read
$\boldsymbol{\hat{\nu}}_{\pm}=([\hat{\lambda}_{\pm}+2(1-zq^{bb}_{s})]/\hat{M}_{ba},1)^{\rm T}.$ (S55)
Since all matrix entries in (S47) are real, the characteristic polynomial also has real coefficients. Therefore, if the eigenvalues are complex, they come in complex-conjugate pairs. The perturbation $\mathbf{\delta m}(t)$ grows in time when ${\rm Re}(\hat{\lambda}_{\pm})>0$, and shrinks in time when ${\rm Re}(\hat{\lambda}_{\pm})<0$. The imaginary part of the eigenvalues tells us whether the perturbation develops oscillations in time ${\rm Im}(\hat{\lambda}_{\pm})\neq 0$, or is monotonic in time ${\rm Im}(\hat{\lambda}_{\pm})=0$. The Hopf bifurcation, also known as an oscillatory instability or type-${\rm II}_{\rm o}$ instability in the Cross-Hohenberg classification [3], occurs when the complex conjugate eigenvalues cross the imaginary axis in the complex plane. Based on Eq. (S52), this occurs when
${\rm tr}(\mathbf{\hat{M}})$ $=0\rightarrow q^{aa}_{s}+q^{bb}_{s}=2/z,$ (S56)
${\rm det}(\mathbf{\hat{M}})$ $>0\rightarrow(zq^{aa}_{s}-1)(zq^{bb}_{s}-1)-\hat{M}_{ab}\hat{M}_{ba}>0.$ (S57)
For fixed $z$, both equations can be solved explicitly to obtain expressions for the critical values of the parameters $J_{\mu}$ and $K_{\mu}$ on the line of Hopf bifurcations. Note that Eq. (S56) sets a direct constraint on the steady-state local order values.
S4.1.1 Perfectly nonreciprocal setting
Focusing on the perfect nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, we have $q^{aa}_{s}=q^{bb}_{s}\equiv q_{s}(J,K)$ and $\hat{M}_{ab}=-\hat{M}_{ba}$. Under these conditions, Eqs. (S56)-(S57) transform into,
${\rm tr}(\mathbf{\hat{M}})$ $=0\rightarrow q_{s}=1/z,$ (S58)
${\rm det}(\mathbf{\hat{M}})$ $>0\rightarrow\hat{M}_{ab}\neq 0,$ (S59)
where the latter equation is directly satisfied for $K\neq 0$ (a proof is given in S4.2.2.). Hence, also in the “monomer” expression we find a critical value for the local order, which is given by $q_{\rm crit}\equiv 1/z$. Upon inserting Eq. (S48) into Eq. (S58), one can obtain an explicit result for the Hopf line in the $(J,K)-$plane.
S4.2 The Bethe-Guggenheim “pair” approximation
In the BG approximation, we can no longer neglect the local order for the linear stability analysis. Hence, we consider a small perturbation of the global and local order around their respective steady-state values
$\mathbf{m}(t)$ $=\mathbf{m}_{s}+\delta\mathbf{m}(t),$ (S60)
$\mathbf{q}(t)$ $=\mathbf{q}_{s}+\delta\mathbf{q}(t).$ (S61)
For the sake of simplicity, we directly focus on the perfectly nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$.
S4.2.1 Steady-state local order
The trivial steady state is given by the disordered state with $m^{\mu}_{s}=0$ and $q^{ab}_{s}=0$. To solve for the steady state of the local order, denoted as $q^{aa}_{s}=q^{bb}_{s}\equiv q_{s}(J,K)$, we need to solve
$q_{s}(J,K)=\frac{1}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{a}_{l,n})},$ (S62)
where $\overline{\mathcal{P}}^{\pm}_{l}$ are the probabilities (S43)-(S44) evaluated at steady-state values
$\overline{\mathcal{P}}^{\pm}_{l}\equiv\mathcal{C}^{z}_{l}(1\mp q_{s})^{z-l}(1\pm q_{s})^{l}.$ (S63)
Equation (S62) can be solved for specific integer values of $z$. For example, for $z\in\{2,3\}$ we obtain
$q_{s}(J,K)|_{z=2}$ $=\frac{2-\sqrt{4-\left[\sum_{n=\pm}\tanh{(2J+nK)}\right]^{2}}}{\sum_{n=\pm}\tanh{(2J+nK)}},$ (S64)
$q_{s}(J,K)|_{z=3}$ $=\frac{4{-}\sqrt{16{-}\sum_{n=\pm}[\tanh{(J{+}nK)}{-}3\tanh{(3J{+}nK)}]\sum_{n=\pm}[\tanh{(J{+}nK)}{+}\tanh{(3J{+}nK)}]}}{\sum_{n=\pm}[\tanh{(J{+}nK)}{-}3\tanh{(3J{+}nK)}]}.$ (S65)
For $z=4$ the solution can be compactly written as the solution to a quartic equation
$q_{s}(J,K)|_{z=4}=\mathcal{S}(J,K)-(1/2)\sqrt{-4\mathcal{S}(J,K)^{2}+2\mathcal{H}(J,K)+\mathcal{Q}(J,K)/\mathcal{S}(J,K)},$ (S66)
where we have introduced the auxiliary functions
$\mathcal{H}(J,K)$ $\equiv\frac{3\cosh{(4J)}[\cosh{(4J)}+\cosh{(2K)}]}{\sinh^{2}{(2J)}[\cosh{(4J)}-2\sinh^{2}{(K)}]},$ (S67)
$\mathcal{Q}(J,K)$ $\equiv 16\left(\sum\nolimits_{n=\pm}[\tanh{(4J+nK)}-2\tanh{(2J+nK)}]\right)^{-1},$ (S68)
$\mathcal{S}(J,K)$ $\equiv(1/2)\sqrt{(2/3)\mathcal{H}(J,K)+(\mathcal{Q}(J,K)/6)\left(\mathcal{A}(J,K)+\Delta_{0}(J,K)/\mathcal{A}(J,K)\right)},$ (S69)
$\mathcal{A}(J,K)$ $\equiv 2^{-1/3}\left(\Delta_{1}(J,K)+\sqrt{\Delta^{2}_{1}(J,K)-4\Delta^{3}_{0}(J,K)}\right)^{1/3},$ (S70)
$\Delta_{0}(J,K)$ $\equiv-(3/4)\sum\nolimits_{n=\pm}[\tanh{(2J{+}nK)}+\tanh{(4J{+}nK)}]\sum\nolimits_{n=\pm}[\tanh{(2J{+}nK)}-\tanh{(4J{+}nK)}],$ (S71)
$\Delta_{1}(J,K)$ $\equiv\!216\cosh{(2J)}\!\cosh{(2K)}\!\operatorname{sech}{(4J{-}K)}\!\operatorname{sech}{(4J{+}K)}\!\sinh^{3}{(2J)}\sinh^{2}{(K)}[\cosh{(4J)}{+}\cosh{(2K)}]^{-2}.$ (S72)
Upon inserting the steady-state values, the probability can be expanded up to first order in $\delta\mathbf{m}(t)$ and $\delta\mathbf{q}(t)$, which reduces Eqs. (5) in the main Letter to the following linear set of equations
$\tau\frac{{\rm d}}{{\rm d}t}\begin{pmatrix}\delta\mathbf{m}(t)\\ \delta\mathbf{q}(t)\end{pmatrix}=\begin{pmatrix}\mathbf{M}&\mathbf{0}\\ \mathbf{0}&\mathbf{Q}\end{pmatrix}\begin{pmatrix}\delta\mathbf{m}(t)\\ \delta\mathbf{q}(t),\end{pmatrix},$ (S73)
where $\mathbf{M}(J,K)$ is a $2\times 2$ matrix and $\mathbf{Q}(J,K)$ a $3\times 3$ matrix. Due to the diagonal block structure of the linearized equations, we can handle the perturbations for $\delta\mathbf{m}(t)$ and $\delta\mathbf{q}(t)$ separately.
S4.2.2 Linear stability of global order
The elements of the $2\times 2$ matrix $\mathbf{M}$ are given by
$M_{aa}(J,K)$ $=M_{bb}(J,K)=\frac{q_{s}/q_{\rm crit}-1}{1+q_{s}}[1-2\sum_{l=0}^{z}\sum_{n=0}^{1}\overline{\mathcal{P}}^{+}_{l}\tanh(U^{a}_{l,n})],$ (S74)
$M_{ab}(J,K)$ $={-}M_{ba}(J,K)=\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\tanh(U^{a}_{l,n}),$ (S75)
where $q_{\rm crit}\equiv 1/(z-1)$. Note that Eqs. (S74)-(S75) are equivalent to Eqs. (7) in the Letter. Further analysis regarding the solution for $\delta\mathbf{m}(t)$ is given in the Letter. Here, we are left to prove that $M_{ab}(J,K\neq 0)\neq 0$. To see this, let us explicitly write out the first sum over $n$
$M_{ab}=\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\tanh(U^{a}_{l,n})=\sum_{l=0}^{z}(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})\left(\tanh{([2l{-}z]J{+}K)}{-}\tanh{([2l{-}z]J{-}K)}\right),$ (S76)
Note that $\overline{\mathcal{P}}^{\pm}_{l}\geq 0$ for $q_{s}\in[-1,1]$, which follows straightforwardly from Eq. (S63). Furthermore, since $\tanh{(x)}$ is an increasing function of $x$, we have $\tanh{([2l-z]J+K)}-\tanh{([2l-z]J-K)}>0$ for $K>0$ and $\tanh{([2l-z]J+K)}-\tanh{([2l-z]J-K)}<0$ for $K<0$. Hence, $M_{ab}$ is given by a sum over strictly positive (for $K>0$) or negative (for $K<0$) terms, rendering $M_{ab}\neq 0$ for $K\neq 0$. This outcome results in complex eigenvalues for $\mathbf{M}$, as explained in the Letter.
S4.2.3 Linear stability of local order
The elements of the $3\times 3$ matrix $\mathbf{Q}$ are given by
$Q_{11}$ $=Q_{22}=-2-\frac{2zq^{2}_{s}}{1-q^{2}_{s}}+\frac{2/z}{1-q^{2}_{s}}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)^{2}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{a}_{l,n})},$ (S77)
$Q_{12}$ $=Q_{21}=0,$ (S78)
$Q_{13}$ $=-Q_{23}=\frac{2}{z}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)(2n-1)(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{a}_{l,n})},$ (S79)
$Q_{31}$ $=-Q_{32}=\frac{1}{1-q^{2}_{s}}\sum_{l=0}^{z}\sum_{n=0}^{1}(2n-1)[(2l-z)(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})-zq_{s}(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})]\tanh{(U^{a}_{l,n})},$ (S80)
$Q_{33}$ $=-2+\sum_{\mu}\sum_{l=0}^{z}\sum_{n=0}^{1}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{\mu}_{l,n})}.$ (S81)
The solution of the linearized equation for $\delta\mathbf{q}(t)$ reads
$\delta\mathbf{q}(t)=\sum_{i=1}^{3}\mathcal{A}_{i}\mathrm{e}^{\tilde{\lambda}_{i}t/\tau}\boldsymbol{\tilde{\nu}}_{i},$ (S82)
where the $\mathcal{A}_{i}$ are determined by the initial conditions, and the eigenvalues of $\mathbf{Q}$, denoted as $\tilde{\lambda}_{i}$, are given by
$\tilde{\lambda}_{1}$ $=Q_{11},$ (S83)
$\tilde{\lambda}_{2}$ $=\frac{1}{2}\left(Q_{11}+Q_{33}+\sqrt{8Q_{13}Q_{31}+(Q_{11}-Q_{33})^{2}}\right),$ (S84)
$\tilde{\lambda}_{3}$ $=\frac{1}{2}\left(Q_{11}+Q_{33}-\sqrt{8Q_{13}Q_{31}+(Q_{11}-Q_{33})^{2}}\right),$ (S85)
and finally, the eigenvectors of $\mathbf{Q}$, denoted as $\boldsymbol{\tilde{\nu}}_{i}$, read
$\boldsymbol{\tilde{\nu}}_{1}$ $=(1,1,0)^{\rm T},$ (S86)
$\boldsymbol{\tilde{\nu}}_{2}$ $=(-Q_{13}/(\tilde{\lambda}_{2}-Q_{33}),Q_{13}/(\tilde{\lambda}_{2}-Q_{33}),1)^{\rm T},$ (S87)
$\boldsymbol{\tilde{\nu}}_{3}$ $=(-Q_{13}/(\tilde{\lambda}_{3}-Q_{33}),Q_{13}/(\tilde{\lambda}_{3}-Q_{33}),1)^{\rm T}.$ (S88)
Note that $\tilde{\lambda}_{1}$ is always real, and therefore cannot give rise to a Hopf bifurcation. The second and third eigenvalues can become complex, in which case their real part is given by ${\rm Re}(\tilde{\lambda}_{2,3})=(Q_{11}+Q_{33})/2$. However, we now prove that $Q_{11}\leq 0$ and $Q_{33}\leq 0$, and therefore also $\tilde{\lambda}_{2,3}$ cannot give rise to a Hopf bifurcation.
To prove that $Q_{11}\leq 0$ we proceed with the following chain of inequalities
$Q_{11}$ $=-2-\frac{2zq^{2}_{s}}{1-q^{2}_{s}}+\frac{2/z}{1-q^{2}_{s}}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)^{2}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{a}_{l,n})}$
$\leq-2-\frac{2zq^{2}_{s}}{1-q^{2}_{s}}+\Bigl{|}\frac{2/z}{1-q^{2}_{s}}\sum_{l=0}^{z}\sum_{n=0}^{1}(2l-z)^{2}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{a}_{l,n})}\Bigr{|}$
$\leq-2-\frac{2zq^{2}_{s}}{1-q^{2}_{s}}+\frac{2/z}{1-q^{2}_{s}}\sum_{l=0}^{z}\sum_{n=0}^{1}|(2l-z)^{2}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{a}_{l,n})}|$
$\leq-2-\frac{2zq^{2}_{s}}{1-q^{2}_{s}}+\frac{2/z}{1-q^{2}_{s}}\sum_{l=0}^{z}|(2l-z)^{2}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})||\tanh{([2l-z]J+K)}+\tanh{([2l-z]J-K)}|$
$\leq-2-\frac{2zq^{2}_{s}}{1-q^{2}_{s}}+\frac{4/z}{1-q^{2}_{s}}\sum_{l=0}^{z}(2l-z)^{2}|\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l}|$
$\leq-2-\frac{2zq^{2}_{s}}{1-q^{2}_{s}}+\frac{4/z}{1-q^{2}_{s}}\sum_{l=0}^{z}(2l-z)^{2}(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})=0,$ (S89)
where the last inequality follows from the triangle inequality $|\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l}|\leq|\overline{\mathcal{P}}^{+}_{l}|+|\overline{\mathcal{P}}^{-}_{l}|$ together with $\overline{\mathcal{P}}^{\pm}_{l}\geq 0$ and therefore $|\overline{\mathcal{P}}^{\pm}_{l}|=\overline{\mathcal{P}}^{\pm}_{l}$. Next, we proceed with $Q_{33}$ in a similar fashion
$Q_{33}$ $=-2+\sum_{\mu}\sum_{l=0}^{z}\sum_{n=0}^{1}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\tanh{(U^{\mu}_{l,n})}$
$=-2+2\sum_{l=0}^{z}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\left[\tanh{([2l-z]J+K)}+\tanh{([2l-z]J-K)}\right]$
$\leq-2+\Bigl{|}2\sum_{l=0}^{z}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\left[\tanh{([2l-z]J+K)}+\tanh{([2l-z]J-K)}\right]\Bigr{|}$
$\leq-2+2\sum_{l=0}^{z}\Bigl{|}(\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l})\left[\tanh{([2l-z]J+K)}+\tanh{([2l-z]J-K)}\right]\Bigr{|}$
$\leq-2+4\sum_{l=0}^{z}|\overline{\mathcal{P}}^{+}_{l}-\overline{\mathcal{P}}^{-}_{l}|$
$\leq-2+4\sum_{l=0}^{z}(\overline{\mathcal{P}}^{+}_{l}+\overline{\mathcal{P}}^{-}_{l})=-2+2=0,$ (S90)
where for the last inequality we again used the triangle inequality. This establishes that $Q_{11}\leq 0$ and $Q_{33}\leq 0$, and therefore when $\tilde{\lambda}_{2,3}$ become complex, their real part obeys the bound ${\rm Re}(\tilde{\lambda}_{2,3})=(Q_{11}+Q_{33})/2\leq 0$. Hence, up to first order, any perturbation $\delta\mathbf{q}(t)$ decays over time. The coherent oscillations in $\mathbf{q}(t)$ observed in Fig. 1c,d in the main Letter are therefore an inherently nonlinear effect related to the coupling between $\mathbf{m}(t)$ and $\mathbf{q}(t)$.
Appendix S5 Divergence of the residence time upon approaching the SNIPER bifurcation
Consider the perfectly nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$. Figure 2c in the main Letter shows the magnetization oscillating between $4$ ghost states upon approaching the degenerate saddle-node-infinite-period (SNIPER) bifurcation from below, i.e. $J\uparrow J_{\rm SNP}(K)$, where $J_{\rm SNP}(K)$ is the $J$ value of the SNIPER bifurcation for a given nonreciprocal coupling $K$. The residence time $\mathcal{T}$ in the vicinity of these ghost states diverges at the SNIPER bifurcation as (see also Eq. (9) in the main Letter)
$\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2},$ (S91)
which is shown explicitly in Fig. S1. To see the precise algebraic scaling, we plot $1/\mathcal{T}^{2}$ against $J_{\rm SNP}(K)-J$ in Fig. S1b, where we indeed observe very good agreement with the expected scaling (black dashed line) close to the SNIPER bifurcation.
Figure S1: Algebraic divergence of the residence time $\mathcal{T}$ within a ghost state close to the degenerate saddle-node-infinite-period (SNIPER) bifurcation. Approaching the SNIPER bifurcation from below, the residence time scales as $\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2}$, where $J_{\rm SNP}(K)$ is the $J$ value of the SNIPER bifurcation for a given nonreciprocal coupling $K$. In both panels we consider $K=0.1$ and $z=4$, for which $J_{\rm SNP}\approx 0.4477$. Results are obtained with the numerical continuation package MATCONT [4].
Appendix S6 Proof of existence of static global order on the finite square lattice
Here, we show that for any finite system size ($N<\infty$) there is at least a regime with $0<|K|\leq\mathcal{O}(1/N)$ where the static global order in the two-dimensional square-lattice is not destroyed. Moreover, in the thermodynamic scaling limit $N\rightarrow\infty$ and $|K|N={\rm constant}>0$ we show that there exists a regime for a nonzero magnetic field $h>0$ and for any coupling strength $J\geq J_{0}$ where the static global order must be preserved.
S6.1 Recap of the global order
[…]
$m^{\mu}(t)\equiv N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}P(\boldsymbol{\sigma};t),$ (S92)
where $\{\boldsymbol{\sigma}\}$ denotes the set of all possible spin configurations $\boldsymbol{\sigma}$. From the master equation (S2) we can obtain the time-evolution equation for the magnetization [2]
$\frac{{\rm d}m^{\mu}(t)}{{\rm d}t}=-2N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}w^{\mu}_{i}(\sigma^{\mu}_{i})P(\boldsymbol{\sigma};t).$ (S93)
In the steady-state we have ${\rm d}m^{\mu}(t)/{\rm d}t=0$, and therefore
$\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}w^{\mu}_{i}(\sigma^{\mu}_{i})P_{\rm s}(\boldsymbol{\sigma})=0,$ (S94)
where $P_{\rm s}(\boldsymbol{\sigma})\equiv\lim_{t\rightarrow\infty}P(\boldsymbol{\sigma};t)$ denotes the steady-state probability.
S6.2 Steady-state for $K=0$
Let us first recall the steady-state global order for $K=0$, where we have two independent Ising systems on lattice $a$ and $b$, respectively. Let $P_{0,\rm s}(\boldsymbol{\sigma})\equiv\lim\limits_{t\rightarrow\infty}P(\boldsymbol{\sigma};t)|_{K=0}$ be the steady-state probability for $K=0$ and an arbitrary magnetic field $h$ (which we need later), and
$w_{0,i}^{\mu}(\sigma^{\mu}_{i})\equiv(1/2\tau)\left[1-\sigma^{\mu}_{i}\tanh{\left(J\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}+h\right)}\right]$ (S95)
the transition rate when $K=0$ in the presence of a magnetic field $h$. The steady-state global order for $K=0$, denoted as $m^{\mu}_{0,\rm s}$, reads [5]
$m^{\mu}_{0,\rm s}\equiv\lim\limits_{h\rightarrow 0^{\pm}}N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}P_{0,\rm s}(\boldsymbol{\sigma}),$ (S96)
which is nonzero for $J\gtrsim\ln{(1+\sqrt{2})}/2$ when $N\gg 1$ (for $N\rightarrow\infty$ the sign $\gtrsim$ changes to $>$). Note that the limit $h\rightarrow 0^{\pm}$ of the magnetic field weakly breaks the $\mathbb{Z}_{2}$-symmetry, such that the steady-state magnetization does not correspond to the unstable value with $m^{\mu}_{0,s}=0$.
S6.3 Perturbation expansion in $K$
We consider a perturbation of the steady-state under a small change in $K$ such that $|K|\ll 1$. For small perturbations we can expand the transition rates as follows,
$w_{i}^{\mu}(\sigma^{\mu}_{i})=w_{0,i}^{\mu}(\sigma^{\mu}_{i})+K\delta w_{i}^{\mu}(\sigma^{\mu}_{i})+\mathcal{O}(K^{2}),$ (S97)
where it follows from a Taylor expansion of Eq. (S3) that
$\delta w_{i}^{\mu}(\sigma^{\mu}_{i})=(1/2\tau)[1-2\delta_{\mu,b}]\sigma^{a}_{i}\sigma^{b}_{i}\operatorname{sech}^{2}{\left(J\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}+h\right)},$ (S98)
with $\delta_{\mu,b}=1$ when $\mu=b$ and $\delta_{\mu,b}=0$ when $\mu=a$. A small perturbation in the transition rates induces a perturbation in the steady-state probability
$P_{\rm s}(\boldsymbol{\sigma})=P_{0,\rm s}(\boldsymbol{\sigma})+K\delta P_{\rm s}(\boldsymbol{\sigma})+\mathcal{O}(K^{2}),$ (S99)
which in turn results in a perturbation of the magnetization, $m^{\mu}_{\rm s}=m^{\mu}_{0,\rm s}+K\delta m^{\mu}_{\rm s}+\mathcal{O}(K^{2})$, where it follows from Eq. (S96) that
$\delta m^{\mu}_{\rm s}\equiv\lim\limits_{h\rightarrow 0^{\pm}}N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma}).$ (S100)
Our aim is to provide an upper and lower bound for $\delta m^{\mu}_{\rm s}$.
S6.4 Bound on perturbations of global order
We start with an upper bound for $\delta m^{\mu}_{\rm s}$, which goes as follows:
$\delta m^{\mu}_{\rm s}$ $\leq\lim\limits_{h\rightarrow 0^{\pm}}|N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$\leq\lim\limits_{h\rightarrow 0^{\pm}}N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=\lim\limits_{h\rightarrow 0^{\pm}}N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}||\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=\lim\limits_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|.$ (S101)
In exactly the same way, we can also provide a lower bound
$\delta m^{\mu}_{\rm s}$ $\geq\lim\limits_{h\rightarrow 0^{\pm}}-|N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$\geq\lim\limits_{h\rightarrow 0^{\pm}}-N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=\lim\limits_{h\rightarrow 0^{\pm}}-N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}||\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=\lim\limits_{h\rightarrow 0^{\pm}}-\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$ (S102)
Hence, combining the upper and lower bound, we obtain
$|\delta m^{\mu}_{\rm s}|\leq\lim\limits_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|.$ (S103)
This is our first main result. We are left with determining an upper bound for $\lim_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, for which we use the following theorem shown in [6] (Theorem 2.1):
Theorem 1
Let the Markov chain $X(t)$ with infinitesimal generator $\mathbf{A}$, i.e., ${\rm d}\mathbf{p}(t)/{\rm d}t=\mathbf{p}(t)\mathbf{A}$, be exponentially weakly ergodic; that is, for any normalized initial conditions $\mathbf{p}(t=0)$, and $\mathbf{p}^{\dagger}(t=0)$, and any $t\geq 0$, there exists a $b>0$ and $c>2$ such that
$\|\mathbf{p}(t)-\mathbf{p}^{\dagger}(t)\|\leq c{\rm e}^{-bt},\ t\geq 0,$ (S104)
where $\|\mathbf{p}\|=\sum_{i}|p_{i}|$ denotes the $l_{1}$-norm for vectors. Then, for perturbations to the infinitesimal generator, $\mathbf{A}+\mathbf{\hat{A}}$, the following bound takes place for the perturbed stationary probabilities $\mathbf{\hat{p}}_{\rm s}$:
$\|\mathbf{p}_{\rm s}-\mathbf{\hat{p}}_{\rm s}\|\leq\frac{1+\ln{(c/2)}}{b}\|\mathbf{\hat{A}}\|,$ (S105)
where $\|\mathbf{\hat{A}}\|=\max_{i}\sum_{j}|\hat{A}_{ij}|$ is the subordinate norm for the perturbation matrix.
Since it is known that the finite volume Glauber dynamics on $\mathbb{Z}^{d}$ is exponentially weakly ergodic (see Theorem 3.3 on page 117 in [7] together with proposition 3.9 on page 124 or simply use Eq. (3.15) in [7]), we can directly use Theorem 1. To translate the results from Theorem 1 into a bound for $|\delta m^{\mu}_{\rm s}|$, we first want to bound the subordinate norm for the perturbed infinitesimal generator. To do so, note that the perturbation to the transition rates obeys the following bound,
$|K\delta w_{i}^{\mu}(\sigma^{\mu}_{i})|=|(K/2\tau)[1-2\delta_{\mu,b}]\sigma^{a}_{i}\sigma^{b}_{i}\operatorname{sech}^{2}{(J\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}+h)}|\leq|K|/2\tau,$ (S106)
where we used that $|\operatorname{sech}^{2}(x)|\leq 1$ for $x\in\mathbb{R}$. These perturbations enter the off-diagonal terms of $\mathbf{\hat{A}}$, so we have established that $\hat{A}_{ij}\leq|K|/2\tau$ for $i\neq j$. Since we consider single spin-flip dynamics, each row/column in $\mathbf{\hat{A}}$ has $2N$ nonzero entries excluding the diagonal entry, which is equal to $\hat{A}_{ii}=-\sum_{j}\hat{A}_{ij}$. Therefore, the diagonal term can also be bounded by $|\hat{A}_{ii}|\leq 2N\times|K|/2\tau$. Combining these results, we obtain the following bound for the subordinate norm of the perturbed infinitesimal generator,
$\|\mathbf{\hat{A}}\|\leq 2\times(2N)\times|K|/2\tau=2N|K|/\tau.$ (S107)
Finally, we note that $\|\mathbf{p}_{\rm s}-\mathbf{\hat{p}}_{\rm s}\|$ in Theorem 1 translates into $\lim_{h\rightarrow 0^{\pm}}|K|\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$ in our work. Combining all together, we obtain that there exists $b>0$ and $c>2$ such that for $K\neq 0$
$\lim\limits_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|\leq\frac{2N(1+\ln{(c/2)})}{b\tau},$ (S108)
and therefore the perturbation is bounded by a finite number (when $N<\infty$ and $\tau\neq 0$) independent of $K$ up to first order
$|\delta m^{\mu}_{\rm s}|\leq\frac{2N(1+\ln{(c/2))}}{b\tau}<\infty.$ (S109)
S6.5 Connecting the dots
We have shown that up to first order the static global order can be written as $m^{\mu}_{\rm s}=m^{\mu}_{0,\rm s}+K\delta m^{\mu}_{\rm s}+\mathcal{O}(K^{2})$ under a small perturbation in $K$, where $m^{\mu}_{0,\rm s}\neq 0$ for $J\gtrsim\ln{(1+\sqrt{2})}/2$ as shown in Eq. (S96). Furthermore, the perturbation $\delta m^{\mu}_{\rm s}$ is bounded by a finite number independent of $K$, as shown in Eq. (S109). This means that for arbitrarily small nonzero $K$, the perturbed global order gets arbitrary close to the unperturbed value, i.e.
$|m^{\mu}_{\rm s}-m^{\mu}_{0,\rm s}|\leq|K|\left(\frac{2N(1+\ln{(c/2)})}{b\tau}\right).$ (S110)
To be more precise, if we set
$0<|K|\leq\mathcal{O}(1/N),$ (S111)
the perturbed steady-state magnetization must come arbitrarily close to the unperturbed steady-state magnetization by merely changing the proportionality factor. If the static global order vanished for any arbitrary nonzero $K$ on the square lattice, it would indicate that $m^{\mu}_{\rm s}$ is a discontinuous and non-differentiable function of $K$ for $J\gtrsim\ln{(1+\sqrt{2})}/2$, as it would suddenly jump from $m^{\mu}_{0,\rm s}\neq 0$ to $m^{\mu}_{\rm s}=0$ in the limit $K\rightarrow 0^{+}$. This cannot be true, since we have established that $m^{\mu}_{\rm s}$ gets arbitrarily close to $m^{\mu}_{0,\rm s}$ up to first order in $K$. Hence, the static global order is not destroyed for finite system sizes ($N<\infty$) in the perfectly nonreciprocal Ising model at least when $0<K\leq\mathcal{O}(1/N)$.
Note that our proof does not state anything about the regime with coherent steady-state oscillations on the square lattice. Similar to the results in [1], we also observe that spiral defects can destroy coherent steady-state oscillations on the square lattice. This, however, does not have any implication for the regime with static order, which is our main focus in this Section.
S6.6 What happens in the thermodynamic limit?
What happens to the static global order when we take $N\rightarrow\infty$? It is known that the infinite volume Glauber dynamics is not ergodic as the Gibbs measure is not unique [7]. Therefore, one cannot use Theorem 1. However, in the presence of a magnetic field with $h>0$ there exists a region $J\geq J_{0}$ (this includes the infinite $J\rightarrow\infty$ limit) such that the infinite volume Glauber dynamics has exponential convergence, i.e., is weakly ergodic, as shown in Theorem 5.1b in [8] (page 479). Hence, in this regime we can use Theorem 1 to obtain a bound for $|K|\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, which would also be given by Eq. (S108) (albeit that $b$ and $c$ might change due to the presence of $h>0$). Looking at Eq. (S110) we can then take a scaling limit $N\rightarrow\infty$ and $|K|N={\rm constant}$, such that the perturbed steady-state long-range order can get arbitrarily close to the unperturbed value by simply changing the proportionality constant. Hence, we conclude that there exists a regime for $|K|N={\rm constant}$ where the static order is preserved when
$N\xrightarrow[h>0,\ J\geq J_{0}]{|K|N={\rm const.}}\infty.$ (S112)
Appendix S7 Kinetic Monte-Carlo simulations
For the results shown in Fig. 1c,d (black dashed lines) and Fig. 3 in the Letter we performed kinetic MC simulations on three different types of lattices: (i) the fully-connected mean field (MF) lattice, (ii) the Bethe lattice, and (iii) the square lattice with periodic boundary conditions. Simulations on the Bethe lattice were performed using the random graph algorithm [9, 10], which works as follows: Let us consider a Bethe lattice with coordination number $z$. First, we create a Cayley tree of $i=\{1,...,N\}$ spins with coordination number $z$. The spins on the outer layer are only connected to one spin on the inner layer. To create the remaining $z-1$ connections, we randomly pair spins on the outer layer to other spins on the outer layer. The final result is a Cayley tree with random connections on the outer layer. Note that for both lattices $a$ and $b$ we create new random connections. For large $N$ it has been shown that the Ising model on an ensemble of such random graphs is equivalent to the Ising model on a Bethe lattice [9]. Indeed, for large $N$ we find perfect agreement between the simulations and our theory, as shown in Fig. 1c,d in the main Letter. For the MF lattice, we connect all spins with each other, resulting in a fully connected graph.
Figure S2: Oscillations in the global order $m^{\mu}(t)$ on the mean-field lattice (a), Bethe lattice (b), and square lattice (c), obtained by Monte-Carlo simulations. In each panel, we plot 10 independent trajectories for the magnetization on lattice $a$ (green) and $b$ (blue). Fluctuations between the trajectories lead to small differences over time, which are decreasing with increasing system size, a signature of self-averaging. The values for $J$ and $K$ for each of the lattices are reported in table S1.
S7.1 Simulation setup
In Table S1 we summarize the system size, number of trajectories, and parameter settings, which were used to obtain the spectral density shown in Fig. 3 in the Letter. As initial conditions, we selected a randomly mixed configuration of up and down spins for fixed magnetization. A subset of trajectories for the global order are displayed in Fig. S2, and in each of the lattices we observe coherent oscillations.
Table S1: Simulation parameters for results shown in Fig. 3 in the main Letter.
lattice size ($N$) MC steps # trajectories $J$ $[k_{\rm B}T]$ $K$ $[k_{\rm B}T]$
mean field $1500$ $1500\times 10^{3}$ $500$ $1.5/N$ $0.3$
Bethe $1457$ $1457\times 10^{3}$ $500$ $0.5$ $0.3$
square $40\times 40$ $1600\times 10^{3}$ $500$ $0.6$ $0.3$
To obtain the results shown in Fig. 1c,d we used a Bethe lattice with system size $N=118097$. Such a large system size was not feasible for the setup of Fig. 3 since the spectral density $\langle|\hat{\Psi}^{\mu}_{k}(\omega)|\rangle$ must be averaged over many independent trajectories, resulting in memory issues for too large $N$.
S7.2 Videos
To visualize the dynamics of the nonreciprocal Ising model on the mean-field lattice, Bethe lattice, and square lattice, we have provided videos as Supplementary Material. In each of the videos, we show three independent simulations on the aforementioned lattices, with the coupling strengths $(J,K)$ reported in Table S1. Further information about the videos is given below:
  • In the video “MF_lattice.gif” simulations on the mean-field lattice are shown for $N=2000$ spins, where each spin is connected to every other spin. For illustrative purposes, the edges between spins are not shown.
  • In the video “Bethe_lattice.gif” simulations on the Bethe lattice are shown for $N=13121$ spins, corresponding to a Bethe lattice with $8$ layers and a coordination number of $z=4$. For illustrative purposes, only the first $6$ layers of the Bethe lattice are shown.
  • In the video “Square_lattice.gif” simulations on the square lattice are shown for $N=122\times 122$ spins.
S7.3 Eigenvalues of Laplacian matrix
Figure S3: Eigenvalues $l_{k}$ of the Laplacian matrix $\mathbf{L}$ for the mean-field lattice (a), Bethe lattice (b), and square lattice (c). For the mean-field lattice, all eigenvalues except for the first are degenerate with value $l_{2}=...=l_{N}=N$, where $N$ is the system size given in table S1. The Bethe lattice has a spectral gap between the lowest eigenvalue $l_{1}$ and $l_{2}$.
In Fig. 3 in the main Letter we plot the spectral density $\langle|\hat{\Psi}^{\mu}_{k}(\omega)|\rangle$ as a function of the eigenvalues $l_{k}$ of the Laplacian matrix $\mathbf{L}$. To obtain the eigenvalues, we numerically diagonalized the Laplacian $\mathbf{L}$ in Python, and the resulting eigenvalues are shown in Fig. S3. Note that for the mean-field lattices all eigenvalues except for the first are degenerate and equal to $l_{2}=..=l_{N}=N$.